OpenStudy (anonymous):

How would I solve the improper integral ʃ xe¯ᵡ² from 0 to infinity

OpenStudy (anonymous):

thats improper Integral

OpenStudy (anonymous):

This is an innocent enough looking integral. However, because infinity is not a real number we can’t just integrate as normal and then “plug in” the infinity to get an answer.

OpenStudy (anonymous):

I understand what you mean the integral isn't the trouble its the bounds that's messing with me

OpenStudy (anonymous):

When you solve this it comes to - e¯ᵡ² / 2 but how would I evaluate

OpenStudy (anonymous):

Length of wire = 10 cm. Length of Side of Square = s Length of Side of Triangle = t Area of Square = A(s) Area of Triangle = A(t) Perimeter of Square = Ps Perimeter of Triangle = Pt Total Area = A(x) Ps = 4x Pt = 10 - Ps Pt = 10 - 4x t = 1/3 Pt t = 1/3 (10 - 4x) A(x) = x² A(t) = t² (√3) / 4 A(t) = 0.443 t² A(t) = 0.433 [1/3 (10 - 4x)]² A(x) = x² + 0.433 [1/3 (10 - 4x)]² A(x) = x² + 0.433 (10/3 - 4/3 x)² A(x) = x² + 0.443 [100/9 - (40/9 x) - (40/9 x) + 16x²/9] A(x) = x² + 0.443 (100/9 - 80/9 x + 16x²/9) A(x) = x² + 0.443 (11.11 - 8.88x + 1.78x²) A(x) = x² + 0.443 (1.78x² - 8.88x + 11.11) A(x) = x² + 0.79x² - 3.9x + 4.9 A(x) = 1.79x² - 3.9x + 4.9 Converting to Standard Form, A(x) = (1.79x² - 3.9x) + 4.9 A(x) = 1.79(x² - 2.11x) + 4.9 A(x) = 1.79(x² - 2.11x +1.11) + 4.9 - 1.79(1.11) A(x) = 1.79(x - 1.06)² + 4.9 - 2 A(x) = 1.79(x - 1.06)² + 2.9 ¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯ Diagram: http://i533.photobucket.com/albums/ee339

OpenStudy (anonymous):

try \[u=-x^2, du=-2xdx\]

OpenStudy (anonymous):

\[-\frac{1}{2}e^{-x^2}\]is the anti derivative, take the limit as x goes to infinity, get 0, replace x by 0 and get \(-\frac{1}{2}\) so your integrals is \(0-(-\frac{1}{2})=\frac{1}{2}\)

OpenStudy (anonymous):

First write x^2 as t then our equation reduces to \[\int\limits_{0}^{\infty}e^{-t}t^{1/2}dt\] limit is again from 0 to infinity which is obvious Now this problem is reduced to taking a Laplace Transform of t^(1/2) for the value of s=1 Hence the solution is \[\sqrt{pi}/2\]

OpenStudy (anonymous):

Thre is no way you can solve this problem without having an idea of laplace transform. laplace transform of the function f(t) is \[\int\limits_{0}^{\infty}e^{-st}f(t)dt\] where s is a parameter. I used this concept above & stating the fact that Laplace transform of t^(1/2) is \[\sqrt{pi}/(2*s^{3/2})\] Value of s is 1 here.